# On The Impact of Replacing Private Cars with Autonomous Shuttles: An Agent-Based Approach

### 23 Nov 2023 v1

@@ Abstract
[…]
@@ I Introduction
Research Gap. Several related works analyze the effect of SAVs on traffic [6, 7, 8, 9]. However, models are often built on outdated data [10, 11, 12, 13, 14], while it has been shown that travel demand forecasts can lead to more realistic scenario analyses [15, 16, 17, 10]. To steer AV adoption, political regulations are posed as a viable solution [18, 19]. However, variations of scenarios are missing to derive an improved understanding of a wider range of possible SAV scenarios. Another line of work examines the impacts on single sustainability indicators, such as traffic-based emissions [20, 21, 22]. Holistic approaches cannot be found, as Table I shows.
@@ II Related Work
Travel Demand Forecast. Often, a timeframe beyond 2030 is considered for AVs [15, 16, 30]. Cokyasar et al. [16] examined the energy consumption and mobility effects of SAE level 4 and 5 AVs in 2025 and 2040 while factoring in changes in overall travel demand. Their analysis considered population growth, shifts in vehicle types, and advancements in technology as key factors influencing travel demand. Hidaka and Shiga [10] consider factors such as shifts in age demographics, reduced travel needs due to aging, rising rates of driver’s license surrender, and regional influences up until 2040. Bansal and Kockelman [15] projected the proportion of AVs in the private light-duty transport sector of the United States. However, these analyses are not agent-based. Microscopic simulations commonly build on outdated inputs, not taking into account any projections into the future [11, 12, 13, 14].
@@
Agent-based (S)AV Simulation. In their analysis based on the Multi-Agent Transport Simulation (MATSim), Liu et al. [8] examined SAV travel in Austin, Texas, using different pricing fares, ranging from 0.5 USD to 1.25 USD per mile of SAV usage. They conclude levels of SAV adoption ranging from 9.2% to 50.9% [8]. Ziemke and Bischoff [13] examine the effects of different SAV regulatory scenarios on accessibility in the MATSim Open Berlin Scenario [31]. They utilize the Demand-Responsive Transport (DRT) extension of MATSim, assuming single-passenger SAV service [13]. For instance, they implement regulatory interventions where a random 10% of agents who previously relied on public transportation are transitioned to SAVs [13]. Using a Munich traffic simulation in MATSim, Zwick et al. [14] examined the impact of SAV introduction on traffic noise. Their simulation incorporates the DRT module and creates a service area that covers almost the entire city [14]. The authors introduce SAVs by replacing all car plans of agents with SAV travel or giving free mode choice [14]. Similarly, Bischoff and Maciejewski [6] conducted their SAV analysis using the MATSim DRT module. While commuting traffic has been incorporated, the SAV fleet handles traffic within Berlin by replacing all private car-based travel plans [6]. We observe a small variety of examined SAV introduction concepts in agent-based simulation, which do not allow for appropriate analyses of potential SAV scenarios and variations that come with it.
@@
Sustainability of AV and SAV travel. For the analysis of sustainability impacts, agent-based simulations are very common [28, 32, 4]. Harper et al. estimated the energy use, emissions, and parking requirements resulting from AVs based on the total vehicle kilometer. Their study found that parking spaces would be freed up while energy use and greenhouse gas emissions would increase by up to 3% [28]. The study considered solely driving-related emissions. Patella et al. [4] used a simulation-based approach to model traffic in Rome and the life cycle impact of AVs. The study focused on quantifying car-related emission generation and energy consumption  [4]. They examined the impact of AVs, including a fully electrified fleet, in their experiments. However, the study only examined a hypothetical 100% electric AV scenario.
@@ V Agent-Based Simulation
By comparing the two scenarios for 2050 without introducing autonomous shuttles with our benchmark SC1.0, we can determine the changed load on the transport infrastructure caused by the increased population. We see a noticeable rise in traffic with increases of up to 12.4% for SC2.0 and 22.7% for SC3.0 for hourly traffic across all links, e.g., streets or public transport lines, until 8:00 pm. We can see that the infrastructure is reaching its limits, especially in the SC3.0 scenario, requiring more intelligent mobility solutions, such as the ones examined in our work.
As we have introduced our new populations, we present how we introduced the shuttles, see Table II, with DRT. We allowed for a maximum fleet size of 100,000 SAVs. Agents can call for specific rides while the shuttle that creates the least workload will be chosen to execute the request. SAVs always stay at the last location of drop off until called again. Alternatively, MATSim can provide hubs for shuttles to return to. However, our experiments showed that this leads to high waiting times and additional vehicle kilometers, which is why we decided to use the default setting. We have analyzed both shuttles with four and six seats. Given the determined fleet sizes by the DRT module, only very few rides showed more than four passengers at a time. Thus, we performed our experiments with shuttles carrying up to four passengers.
@@ VII Evaluation
As we build our sustainability analysis primarily on kilometers driven, we compare our findings to those of others in the field. Patella et al. [4] conclude total traffic kilometer variations to range from a 5.0% decrease in urban areas and 8.0% increase on highways in a 100% AV scenario. The decrease in urban areas can be attributed to a shift from urban to highway-based travel as higher capacity gains were applied to highways compared to urban roads [4]. They find a total increase of vehicle distance of 1% [4]. Their approach is similar to our SC1.3, where we predict an increase of up to 1.3%. However, their study assumes AVs rather than SAVs. Liu et al. [8] analyze the SAV adoption based on variations of SAV pricing fares, as outlined in Chapter II. They derive vehicle distance increases ranging from 9.8% to 15.7%, with SAV adoption ranging from 9.2% to 50.9%, relative to all mode trips (including non-car-based travel) [8]. Increased SAV shares show a reduced increase of vehicles distances in the analysis of Liu et al. [8], similar to our findings. For a 69.4% SAV share, we see an increased travel demand of 0.8%, which goes up to 2.8% for a 38.9% SAV share.
As Bischoff and Maciejewski [6] have conducted a SAV analysis for Berlin City, we compare their 100% scenario statistics with key figures of our 100% SAV scenarios in the following. Bischoff and Maciejewski [6] conclude the average SAV to drive 274 kilometers (compared to our 203-216 kilometers in our scenarios), with a 13% empty ride share (compared to our 3% to 3.1%). The average trip distance was described with 9.1 kilometers (compared to our 11.0 kilometers) [6]. The differences can be attributed to Bischoff and Maciejewski [6] focusing on the traffic affecting the Berlin city area, excluding any travel that is not entering or exiting the city (around 35% of the total trips). Additionally, Bischoff and Maciejewski [6] consider a fleet size of 100,000 SAVs for a 100% Open Berlin Scenario. In comparison, we see up to 68,100 SAVs in our fleets for the 10% Open Berlin Scenario. This is due to the inclusion of the rural areas, which require a significantly larger fleet. As a side-effect, we observe decreased average distances traveled per SAV, as the fleet is larger. We also see a decrease in empty rides, which is partly due to the fact that SAVs stay in close proximity to the agent’s home addresses during nighttime.
Sustainability Impacts. In our analysis, we have incorporated changes in the powertrain split by contrasting the current powertrain split to a 100% BEV scenario, similar to Patella et al. [4]. We did not account for future improvements of ICEVs. Also, in contrast to Fagnant et al. [32], we do not differentiate between warm and cold-start emissions. Instead, we mainly differentiate between rural and urban traffic, similar to other scholars (e.g.,  [51, 4]). Additionally, BEV emissions strongly depend on the share of renewable sources of the electricity mix [52]. While we assume to have a 100% renewable electricity mix by 2050, the current electricity mix of Germany equals 41.1% in electricity production and 19.7% in electricity brutto consumption in 2021 [53].
@@ VIII Conclusion
Outlook. While we provided two diverse forecasts of travel demands, the granularity of the approach can be further increased, e.g., by incorporating social changes. As we predicted travel demand, but no changes in traffic infrastructure, e.g. increased road capacity or changes in public transport, a future Open Berlin Scenario for 2050 would be of interest, also incorporating increased travel demands in earlier stages. Generally, increasing the number of iterations for our simulations would increase the accuracy of our results. Waiting times of agents could be partially improved by a scheduling algorithm [54] on the day ahead, as there is no need for real-time dispatching for regular commutes. An additional mode of privately owned AVs would further improve the diversity of our experiments. Other approaches to SAV adoption, especially more naturally evolving ones, could be compared to our scenarios. Finally, improved calculations for our sustainability impacts are of interest. Here, we would like to incorporate cold and warm starts and also the effects of large-scale remote assistance centers, which might be required for large SAV fleets [55, 56]. Here, especially the dimensionality of such centers [57] and their sustainability implications are of interest.

### 19 Jan 2024 v3

@@ Abstract
The European Green Deal aims to achieve climate neutrality by 2050, which demands improved emissions efficiency from the transportation industry. This study uses an agent-based simulation to analyze the sustainability impacts of shared autonomous shuttles. We forecast travel demands for 2050 and simulate regulatory interventions in the form of replacing private cars with a fleet of shared autonomous shuttles in specific areas. We derive driving-related emissions, energy consumption, and non-driving-related emissions to calculate life-cycle emissions. We observe reduced life-cycle emissions from 0.4% to 9.6% and reduced energy consumption from 1.5% to 12.2%.
@@ I Introduction
Research Gap. Several related works analyze the effect of SAVs on traffic [6, 7, 8, 9]. However, models are often built on outdated data [10, 11, 12, 13, 14], while it has been shown that travel demand forecasts can lead to more realistic scenario analyses [15, 16, 17, 10]. To steer AV adoption, political regulations are posed as a viable solution [18, 19]. However, variations of scenarios are missing to derive an improved understanding of a wider range of possible SAV scenarios. Another line of work examines the effects on individual sustainability metrics [20, 21, 22]. Holistic approaches cannot be found, as Table I shows.
@@ II Related Work
Travel Demand Forecast. Often, a timeframe beyond 2030 is considered for AVs [15, 16, 30]. Cokyasar et al. [16] examined AVs’ energy consumption in 2025 and 2040 while factoring in changes in overall travel demand. Their analysis considered population growth, shifts in vehicle types, and advancements in technology as key factors influencing travel demand. Hidaka and Shiga [10] consider factors such as shifts in age demographics, reduced travel needs due to aging, rising rates of driver’s license surrender, and regional influences up until 2040. Bansal and Kockelman [15] projected potential shares of AVs in the private light-duty transport sector of the United States, considering aspects such as pricing and willingness to pay.
However, these analyses are not agent-based. Microscopic simulations commonly build on outdated inputs, not taking into account any projections into the future [11, 12, 13, 14].
@@
Agent-based (S)AV Simulation. In their analysis based on the Multi-Agent Transport Simulation (MATSim), Liu et al. [8] examined the effect of different pricing strategies on SAV travel in Austin, Texas, ranging from 0.5 USD to 1.25 USD per mile. Based on their simulation, they see adoption ranges between 9.2% and 50.9% [8]. Ziemke and Bischoff [13] examine the effects of different SAV regulatory scenarios on accessibility in the MATSim Open Berlin Scenario [31]. They utilize the Demand-Responsive Transport (DRT) extension of MATSim, assuming single-passenger SAV service [13]. For instance, they implement regulatory interventions where a random selection of 10% of agents that were using publication transportation are now being served by SAVs [13], which does not take a holistic view but is an interesting approach. Using a Munich traffic simulation in MATSim, Zwick et al. [14] examined how SAVs affect traffic noise levels on a city-wide level [14]. The authors introduce SAVs by replacing all car plans of agents with SAV travel or giving free mode choice [14]. Similarly, Bischoff and Maciejewski [6] conducted their SAV analysis using the MATSim DRT module. They simulate a scenario where all trips in the city of Berlin based on private cars are being replaced by an SAV fleet [6].
In general, we observe a small variety of examined SAV introduction concepts in agent-based simulation, which do not allow for appropriate analyses of potential SAV scenarios and variations that come with it.
@@
Sustainability of AV and SAV travel. For the analysis of sustainability impacts, agent-based simulations are very common [28, 32, 4]. Harper et al. used traveled vehicle kilometers to calculate emissions and energy consumption for AVs. Their study found that energy use and greenhouse gas emissions would increase by up to 3% [28]. The study considered solely driving-related emissions. Patella et al. [4] performed a simulation of traffic in Rome, including the life cycle impact of AVs. The study focused on quantifying car-related emission generation and energy consumption  [4]. They examined the impact of AVs, including a fully electrified fleet, in their experiments. However, the study only examined a hypothetical 100% electric AV scenario.
@@ V Agent-Based Simulation
By comparing the two scenarios for 2050 without introducing autonomous shuttles with our benchmark SC1.0, we can determine the changed load on the transport infrastructure caused by the increased population. As detailed in [23], we see an increase in traffic of up to 12.4% for SC2.0 and 22.7% for SC3.0 for hourly traffic across all links, e.g., streets or public transport lines, until 8:00 pm. Especially for SC3.0, the infrastructure is reaching its limits, as after-hours traffic causes congestion that lasts well past midnight. Such a case requires more intelligent mobility solutions, such as the ones examined in our work.
As we have introduced our new populations, we present how we introduced the shuttles, see Table II, with DRT. For technical reasons, we allowed for a maximum fleet size of 100,000 SAVs, well above the maximum number of necessary vehicles in any scenario as computed by DRT. Agents can call for specific rides while the shuttle that creates the least workload will be chosen to execute the request. SAVs always stay at the last location of drop off until called again. Alternatively, MATSim can provide hubs for shuttles to return to. However, our experiments showed that this leads to high waiting times and additional vehicle kilometers, which is why we decided to use the default setting. We have analyzed both shuttles with four and six seats. Given the determined fleet sizes by the DRT module, only very few rides showed more than four passengers at a time. Thus, we performed our experiments with shuttles carrying up to four passengers.
@@ VII Evaluation
As we build our sustainability analysis primarily on kilometers driven, we compare our findings to those of others in the field. In a 100% AV scenario, Patella et al. [4] see an increase of 1% for vehicles distances and found decreases of 5.0% in urban areas and increases of 8.0% on highways. They argue that AVs show higher capacity impacts on highways. Their approach is similar to our SC1.3, where we predict an increase of up to 1.3%. However, their study assumes AVs rather than SAVs. Different pricing strategies were examined by Liu et al. [8], where they see shares of shared autonomous vehicles between 9.2% and 50.9%. This leads to increased vehicle distances between 9.8% and 15.7%. Similar to our findings, they see that vehicle distances don’t increase as much anymore when the SAV share is already high. For a 69.4% SAV share, we see an increased travel demand of 0.8%, which goes up to 2.8% for a 38.9% SAV share.
A scenario where 100% of travel in Berlin City was performed by autonomous vehicles was provided by Bischoff and Maciejewski [6]. Our comparable scenarios show that an SAV travels a distance between 203 and 216 km on average, which is similar to their 274 km. Also, trip distances were comparable, with 9.1 km in their experiments and 11.0 km in ours. As we also included traffic leaving and entering the Berlin city area, our number is slightly higher. However, they saw a share of roughly 13% empty rides, compared to our much lower 3%. This can be explained by our much larger fleet size, which became necessary due to the including of outer Berlin. In our experiments, up to 68,100 SAVs were deployed for the 10% Open Berlin Scenario, whereas their experiments see only 100,000 SAVs for the full 100% Open Berlin Scenario. Also, SAVs stay in close proximity to the agent’s home addresses during nighttime.
Sustainability Impacts. Our sustainability analysis includes both driving-related emissions and non-driving related emissions. Neglecting future improvements of ICEVs and differences between warm- and cold-starts [32], we compare a common powertrain split with a fully electrified scenario, as done by Patella et al. [4]. For this, we assume fully renewable electricity, which affects the calculation of emissions drastically [51] and differs strongly from the electricity mix we see today [52]. As we include traffic from both urban and rural areas, we also distinguish between these categories when we compute our sustainability metrics, as previously done [53, 4].
@@ VIII Conclusion
Outlook. While we provided two diverse forecasts of travel demands, the granularity of the approach can be further increased, e.g., by incorporating social changes. As we predicted travel demand but no changes in traffic infrastructure, e.g., increased road capacity or changes in public transport, a future Open Berlin Scenario for 2050 would be of interest, also incorporating increased travel demands in earlier stages. Generally, increasing the number of iterations for our simulations would increase the accuracy of our results. Waiting times of agents could be partially improved by a scheduling algorithm [54] on the day ahead, as there is no need for real-time dispatching for regular commutes. An additional mode of privately owned AVs would further improve the diversity of our experiments. Other approaches to SAV adoption, especially more naturally evolving ones, could be compared to our scenarios. Finally, improved calculations for our sustainability impacts are of interest. Here, we would like to incorporate temporal factors that affect energy consumption, such as cold and warm starts, as well as the effects of large-scale remote assistance centers, which are necessary for large SAV fleets [55, 56]. Here, especially the dimensionality of such centers [57] and their sustainability implications are of interest.